OpenStudy (anonymous):

need the lowest common multiples for: 1. 20 and 35 2. 28 and 35 3. 40 and 70 4. 52 and 160 Thanks

OpenStudy (amistre64):

1. start with 35 till you hit something that 20 goes into

OpenStudy (amistre64):

same for the rest really

OpenStudy (anonymous):

1. 140 2. 140 3. 280 4. 2080

OpenStudy (amistre64):

20/35 = 1 4/7 = 1 *4 *4 -------- 28/7 = 4

OpenStudy (amistre64):

35*4 = 140

OpenStudy (anonymous):

We need to find the smallest number that both integers divide evenly. So take the larger of the two and start multiplying it by 2, 3, 4 etc. until we find a number the smaller number divides evenly. take 20 and 35: 2 * 35 = 70 not divisible by 20 3 * 35 = 103 not divisible by 20 4 * 35 = 140 divisible 20 140/20 = 7 LCM of 20 and 35 = 140 A trick that my help is to take each term and factor it. Then ignore common factors and multiply the remainig factors together. example 28 and 35: 28 = 4 x 7 35 = 5 x 7 so try 4 x 5 x 7 = 140 So LCM of 28 and 35 = 140

OpenStudy (amistre64):

3) 40 and 70 40/70 = 1 4/7 = 1 28/7 = 4 70*4 = 280